OpenStudy (anonymous):

solving linear equations 6.94=2.98+2b

OpenStudy (anonymous):

this was the story problem Jake is going to the store to buy one gallon of milk and 2 loaves of bread. The milk cost 2.98 a gallon. Jake's total was 6.94. How much did each loaf of bread cost?

OpenStudy (texaschic101):

you did set the equation up correctly... 1. subtract 2.98 from both sides 2. divide both sides by 2

OpenStudy (texaschic101):

do that, and you should be able to solve the answer for b (the cost of the bread)

OpenStudy (anonymous):

ok then one sec.

OpenStudy (anonymous):

i got 2.08

OpenStudy (texaschic101):

not correct... what is 6.94 - 2.98 ??

OpenStudy (texaschic101):

then divide that by 2 and what do you get ?

OpenStudy (anonymous):

ohhh it was 3.51

OpenStudy (anonymous):

no it was 3.96 2.98-6.94

OpenStudy (texaschic101):

let me show you.. 6.94 = 2.98 + 2b --- subtract 2.98 from both sides 6.94 - 2.98 = 2.98 - 2.98 + 2b --- simplify 3.96 = 2b -- divide both sides by 2 3.96/2 = 2b/2 1.98 = b each loaf of bread costs $ 1.98

OpenStudy (anonymous):

thanks. :)

OpenStudy (texaschic101):

sure thing :)